OpenStudy (anonymous):

ok so Optimization, i just learned this so bear with me. Find the length and width of a rectangle that has an area of 968 square feet and whose perimeter is a minimum. I was getting the ones where i maximized correct, but how does my approach differ when minimizing?

OpenStudy (anonymous):

do your own homework sir!

OpenStudy (anonymous):

eh! i just don't understand why i cant get the minimum problems right :D if someone could please explain, i already know the answer, but i don't understand

OpenStudy (anonymous):

Im not always correct with optimization. So, ill let someone else help you. You should seek help from @jim_thompson5910 or @phi ..they know they're stuff and are really helpful with explanations.

jimthompson5910 (jim_thompson5910):

A = LW 968 = LW 968/W = L L = 968/W ----------------------- P = 2L + 2W P = 2(968/W) + 2W P = 1936/W + 2W P = 1936/W + (2W^2)/W P = (1936 + 2W^2)/W You want to minimize this, so derive it with respect to W. Then set the derivative equal to 0 and solve for W. This will give you the critical values, then you can use either the first or second derivative test to see which critical values are the min/max.

OpenStudy (anonymous):

hmm i still got the same answer.

jimthompson5910 (jim_thompson5910):

what answer are you getting

OpenStudy (anonymous):

well my x value is 22squareroot2

jimthompson5910 (jim_thompson5910):

that's the value of W that yields the smallest perimeter

jimthompson5910 (jim_thompson5910):

use that to find the length

OpenStudy (anonymous):

but apparently the answer is 484ft by 2ft

jimthompson5910 (jim_thompson5910):

those are one set of dimensions for this rectangle, but they don't produce the smallest perimeter

jimthompson5910 (jim_thompson5910):

the perimeter of that 484ft by 2ft rectangle is P = 2L + 2W P = 2*484 + 2*2 P = 968 + 4 P = 972 --------------------------------------------------------------- but the rectangle 242 ft by 4 ft gives you a smaller perimeter (still not the smallest though) P = 2L + 2W P = 2*242 + 2*4 P = 484 + 8 P = 492 so the answer can't be 484ft by 2ft just by this counter-example alone

jimthompson5910 (jim_thompson5910):

so there's clearly a typo somewhere

OpenStudy (anonymous):

:/ ok i'll ask my teacher tomorrow. thanks for all your help!

jimthompson5910 (jim_thompson5910):

yw